OpenStudy (anonymous):

The two prime numbers p and m are between 50 and 60. Their difference is 6. What is their sum? Tell how you found your answer. Please help

OpenStudy (anonymous):

List the numbers between 50 and 60. 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60. Now erase the even numbers since 2 is a factor of them (which makes them composite and not prime). 51, 53, 55, 57, 59. Then erase 51 (17 is a factor since 17 * 3 = 51), 55 (55 = 5 * 11), and 57 (57 = 3 * 19) to be left with this list. 53, 59. These two numbers are prime since their only factors are 1 and the number itself. Also notice how their difference is 59 - 53 = 6. So that part checks out. Therefore, p = 53, m = 59 Sum: p + m = 53 + 59 = 112. Final Answer: 112